# Probabilistic Guarantees to Explicit Constructions: Local Properties of Linear Codes

### v1

@@ 6 Consequences
Property $\mathcal{P}(\rho,\sigma,\ell,L)$ is a reasonable $(L+1)$-LCL property. In particular, we have
$\kappa_{q}(\mathcal{P}(\rho,\sigma,\ell,L))=\log_{q}(\ell+1)^{(L+1)}.$
Furthermore, $T_{\mathcal{P}(\rho,\sigma,\ell,L)}\leq(\ell+1)^{(L+1)}$.
@@
Lastly, we prove that $\mathcal{P}$ is a reasonable $(L+1)$-LCL property. This is seen by observing that the number of matrices of the form $\mathbf{M}(K,P)$, as constructed above, is at most $(\ell+1)^{(L+1)}$, which is an upper bound on the number of partitions consisting of at most $\ell$ parts, of every subset of $[L+1]$. Thus, $T_{\mathcal{P}}\leq(\ell+1)^{(L+1)}$, and the number of local profiles associated with this property is at most
$(\ell+1)^{(L+1)n}=q^{\log_{q}(\ell+1)^{(L+1)}\cdot n},$
and we see that $\kappa_{q}(\mathcal{P})=\log_{q}(\ell+1)^{(L+1)}$ goes to zero as $q\rightarrow\infty$. ∎
@@
For $\mathcal{P}(\rho,\sigma,\ell,L)$, let $\mathcal{C}_{\textrm{in}},\mathcal{C}_{\textrm{out}},G$ be as defined in Theorem˜5.32. Furthermore, let $\mathcal{C}_{\textrm{out}}\subseteq({\mathbb{F}}_{q}^{R_{\textrm{in}}d})^{N}$ be a code with rate $R_{\textrm{out}}=1-\varepsilon$ and distance $\delta_{\textrm{out}}\geq\varepsilon^{3}$. Then, $\mathcal{C}_{\textrm{AEL}}(\mathcal{C}_{\textrm{in}},\mathcal{C}_{\textrm{out}},G)\subseteq{\mathbb{F}}_{Q}^{N}$ is a ${\mathbb{F}}_{q}$-linear code that is $(\rho,\sigma,\ell,L)$-average radius list-recoverable with erasures, has rate $R_{\textrm{AEL}}>R_{\mathcal{P}}-2\varepsilon$, and can be constructed in time ${\mathrm{poly}}(N)$. Additionally, we have $q\leq(\ell+1)^{(L+1)\cdot\frac{8}{\varepsilon}}$, and $Q=q^{d}$, where
@@
By ˜6.8, $\mathcal{P}(\rho,\sigma,\ell,L)$ is a reasonable property. We now turn to set a value for $q$ so that Eq.˜13 is satisfied. For $q=(\ell+1)^{(L+1)\cdot\frac{8}{\varepsilon}}$, we see
@@
Therefore, Corollary˜6.1 implies that $\mathcal{C}_{\textrm{AEL}}$ does not satisfy $\mathcal{P}(\rho,\sigma,\ell,L)$, and thus, is $(\rho,\sigma,\ell,L)$-average radius list-recoverable with erasures.
Finally, Eq.˜21 from Corollary˜6.1, along with the value of $T_{\mathcal{P}(\rho,\sigma,\ell,L)}$ from ˜6.8 implies
@@
We now record several results, corresponding to the special cases defined above. The proofs for them follow from Corollary˜6.9. The first details parameters for list recovery at capacity.
@@
For a fixed input list size $\ell$, and a fixed radius $\rho>0$, let $L_{\rho,\ell}$ be the smallest output list size such that $R_{\mathcal{P}(\rho,0,\ell,L_{\rho,\ell})}\geq 1-\rho$. For $\varepsilon>0$, let $R:=1-\rho-\varepsilon\leq R_{\mathcal{P}(\rho,0,\ell,L_{\rho,\ell})}-\varepsilon$. The code $\mathcal{C}_{\textrm{AEL}}$ as in Corollary˜6.9 is $(1-R-\varepsilon,\ell,L_{\rho,\ell})$-average radius list-recoverable, with rate $R_{\textrm{AEL}}>R_{\mathcal{P}(\rho,0,\ell,L_{\rho,\ell})}-2\varepsilon$, and alphabet size at most $\exp\bigl(\left(L_{\rho,\ell}/\varepsilon\right)^{O(L_{\rho,\ell})}\bigr)$.
@@
For $\varepsilon>0$ and a fixed input list size $\ell$ and rate $R$, let $L_{R,\ell}$ be the smallest output list size such that $R_{\mathcal{P}(0,0,\ell,L_{R,\ell})}-\varepsilon\geq R$. Then, the code $\mathcal{C}_{\textrm{AEL}}$ as in Corollary˜6.9 is $(\ell,L_{R,\ell})$-zero error list recoverable, with rate $R_{\textrm{AEL}}>R_{\mathcal{P}(0,0,\ell,L_{R,\ell})}-2\varepsilon$, and alphabet size at most $\exp\bigl(\left(L_{R,\ell}/\varepsilon\right)^{O(L_{R,\ell})}\bigr)$.
@@
For constants $\varepsilon>0$, $0\leq\sigma\leq 1$, a fixed input list size $\ell$ and rate $R$, let $L_{\sigma,R,\ell}$ be the smallest output list size such that $R_{\mathcal{P}(0,\sigma,\ell,L_{\sigma,R,\ell})}-\varepsilon\geq R$. Then, the code $\mathcal{C}_{\textrm{AEL}}$ as in Corollary˜6.9 is $(\sigma,\ell,L_{\sigma,R,\ell})$-erasure list recoverable, with rate $R_{\textrm{AEL}}>R_{\mathcal{P}(0,\sigma,\ell,L_{\sigma,R,\ell})}-2\varepsilon$, and alphabet size at most $\exp\bigl(\left(L_{\sigma,R,\ell}/\varepsilon\right)^{O(L_{\sigma,R,\ell})}\bigr)$.
@@
For an integer $t\geq 2$, let $\mathcal{C}_{\textrm{AEL}}$ be the code as in Corollary˜6.9 for the property $\mathcal{P}(0,0,t-1,t-1)$. Then, $\mathcal{C}_{\textrm{AEL}}$ is a code of rate $R=R_{\mathcal{P}(0,0,t-1,t-1)}-2\varepsilon$ that is (0,t1,t1)(0,t-1,t-1)-list recoverable. The alphabet size is at most $\exp((t/\varepsilon)^{O(t)})$ and moreover, the codewords of $\mathcal{C}_{\textrm{AEL}}$, when arranged as columns in a $N\times Q^{RN}$ matrix, form a perfect hash matrix.
@@
The list-recoverability of $\mathcal{C}_{\textrm{AEL}}$ follows from Corollary˜6.9. We prove that the set of codewords of a (0,t1,t1)(0,t-1,t-1)-list-recoverable code can be arranged as columns in a matrix to form a perfect hash matrix. Indeed, (0,t1,t1)(0,t-1,t-1)-list recoverability implies that the number of pairwise distinct codewords having at most t1t-1 distinct entries in every row is at most t1t-1. This implies that for any set of $t$ codewords, there exists at least one index on which the $t$ codewords have pairwise distinct entries. ∎
@@ 7 Acknowledgements
We thank Jonathan Mosheiff for providing feedback and comments on a draft of the paper.

### v2

@@ 6.1 List Recovery
Property $\mathcal{P}(\rho,\sigma,\ell,L)$ is an $(L+1)$-LCL property. Additionally, we have
$\kappa_{q}(\mathcal{P}(\rho,\sigma,\ell,L))\leq\log_{q}(\ell+1)^{(L+1)},$
and
$T_{\mathcal{P}(\rho,\sigma,\ell,L)}\leq(\ell+1)^{(L+1)}.$
@@
Lastly, we prove the upper bounds on $\kappa_{q}(\mathcal{P})$ and $T_{\mathcal{P}}$. This is seen by observing that the number of matrices of the form $\mathbf{M}(K,P)$, as constructed above, is at most $(\ell+1)^{(L+1)}$, which is an upper bound on the number of partitions consisting of at most $\ell$ parts, of every subset of $[L+1]$. Thus, $T_{\mathcal{P}}\leq(\ell+1)^{(L+1)}$, and the number of local profiles associated with this property is at most
$(\ell+1)^{(L+1)n}=q^{\log_{q}(\ell+1)^{(L+1)}\cdot n}$
and we see that $\kappa_{q}(\mathcal{P})\leq\log_{q}(\ell+1)^{(L+1)}$. ∎
@@
For $\mathcal{P}(\rho,\sigma,\ell,L)$, let $\mathcal{C}_{\textnormal{in}},\mathcal{C}_{\textnormal{out}},G$ be as defined in Theorem˜5.32. Furthermore, let $\mathcal{C}_{\textnormal{out}}\subseteq({\mathbb{F}}_{q}^{R_{\textnormal{in}}d})^{N}$ be a code with rate $R_{\textnormal{out}}=1-\varepsilon$ and distance $\delta_{\textnormal{out}}\geq\varepsilon^{3}$. Then, $\mathcal{C}_{\textnormal{AEL}}(\mathcal{C}_{\textnormal{in}},\mathcal{C}_{\textnormal{out}},G)\subseteq{\mathbb{F}}_{Q}^{N}$ is a ${\mathbb{F}}_{q}$-linear code that is $(\rho,\sigma,\ell,L)$-average radius list-recoverable with erasures, has rate $R_{\textnormal{AEL}}>R_{\mathcal{P}}-2\varepsilon$, and can be constructed in time ${\mathrm{poly}}(N)$. Additionally, we have $q\leq(\ell+1)^{(L+1)\cdot\frac{8}{\varepsilon}}$, and $Q=q^{d}$, where
@@
We set a value for $q$ so that Eq.˜13 is satisfied. For $q=(\ell+1)^{(L+1)\cdot\frac{8}{\varepsilon}}$, we see
@@
Therefore, Corollary˜6.1 implies that $\mathcal{C}_{\textnormal{AEL}}$ does not satisfy $\mathcal{P}(\rho,\sigma,\ell,L)$, and thus, is $(\rho,\sigma,\ell,L)$-average radius list-recoverable with erasures.
Finally, Eq.˜21 from Corollary˜6.1, along with the value of $T_{\mathcal{P}(\rho,\sigma,\ell,L)}$ from ˜6.9 implies
@@
We now record several results, corresponding to the special cases defined above. The proofs for them follow from Corollary˜6.10. The first details parameters for list recovery at capacity.
@@
For a fixed input list size $\ell$, and a fixed radius $\rho>0$, let $L_{\rho,\ell}$ be the smallest output list size such that $R_{\mathcal{P}(\rho,0,\ell,L_{\rho,\ell})}\geq 1-\rho$. For $\varepsilon>0$, let $R:=1-\rho-\varepsilon\leq R_{\mathcal{P}(\rho,0,\ell,L_{\rho,\ell})}-\varepsilon$. The code $\mathcal{C}_{\textnormal{AEL}}$ as in Corollary˜6.10 is $(1-R-\varepsilon,\ell,L_{\rho,\ell})$-average radius list-recoverable, with rate $R_{\textnormal{AEL}}>R_{\mathcal{P}(\rho,0,\ell,L_{\rho,\ell})}-2\varepsilon$, and alphabet size at most $\exp\bigl(\left(L_{\rho,\ell}/\varepsilon\right)^{O(L_{\rho,\ell})}\bigr)$.
@@
For $\varepsilon>0$ and a fixed input list size $\ell$ and rate $R$, let $L_{R,\ell}$ be the smallest output list size such that $R_{\mathcal{P}(0,0,\ell,L_{R,\ell})}-\varepsilon\geq R$. Then, the code $\mathcal{C}_{\textnormal{AEL}}$ as in Corollary˜6.10 is $(\ell,L_{R,\ell})$-zero error list recoverable, with rate $R_{\textnormal{AEL}}>R_{\mathcal{P}(0,0,\ell,L_{R,\ell})}-2\varepsilon$, and alphabet size at most $\exp\bigl(\left(L_{R,\ell}/\varepsilon\right)^{O(L_{R,\ell})}\bigr)$.
@@
For constants $\varepsilon>0$, $0\leq\sigma\leq 1$, a fixed input list size $\ell$ and rate $R$, let $L_{\sigma,R,\ell}$ be the smallest output list size such that $R_{\mathcal{P}(0,\sigma,\ell,L_{\sigma,R,\ell})}-\varepsilon\geq R$. Then, the code $\mathcal{C}_{\textnormal{AEL}}$ as in Corollary˜6.10 is $(\sigma,\ell,L_{\sigma,R,\ell})$-erasure list recoverable, with rate $R_{\textnormal{AEL}}>R_{\mathcal{P}(0,\sigma,\ell,L_{\sigma,R,\ell})}-2\varepsilon$, and alphabet size at most $\exp\bigl(\left(L_{\sigma,R,\ell}/\varepsilon\right)^{O(L_{\sigma,R,\ell})}\bigr)$.
@@
For an integer $t\geq 2$, let $\mathcal{C}_{\textnormal{AEL}}$ be the code as in Corollary˜6.10 for the property $\mathcal{P}(0,0,t-1,t-1)$. Then, $\mathcal{C}_{\textnormal{AEL}}$ is a code of rate $R=R_{\mathcal{P}(0,0,t-1,t-1)}-2\varepsilon$ that is (0,t1,t1)(0,t-1,t-1)-list recoverable. The alphabet size is at most $\exp((t/\varepsilon)^{O(t)})$ and moreover, the codewords of $\mathcal{C}_{\textnormal{AEL}}$, when arranged as columns in a $N\times Q^{RN}$ matrix, form a perfect hash matrix.
@@
The list-recoverability of $\mathcal{C}_{\textnormal{AEL}}$ follows from Corollary˜6.10. We prove that the set of codewords of a (0,t1,t1)(0,t-1,t-1)-list-recoverable code can be arranged as columns in a matrix to form a perfect hash matrix. Indeed, (0,t1,t1)(0,t-1,t-1)-list recoverability implies that the number of pairwise distinct codewords having at most t1t-1 distinct entries in every row is at most t1t-1. This implies that for any set of $t$ codewords, there exists at least one index on which the $t$ codewords have pairwise distinct entries. ∎
6.2 Efficient Decoding Algorithms
We remark our derandomization of linear codes admit efficient (list) decoding algorithms via the recently developed efficient (list) decoding algorithms for AEL [JMST25, JS25, ST25]. There are two kinds of such decoders: one based on weak-regularity lemmas [JS25, ST25], and another based on the Sum-of-Squares SDP hierarchy [JMST25]. The former runs in near-linear time in the block length, whereas the latter has a larger polynomial running time. The crucial property needed by these AEL decoders is that the expander graph used in the construction be sufficiently expanding, namely, sufficiently small second largest singular value $\lambda$ of their (bi)adjacency matrix.
For the sake of completeness, we recall the main decoding results of [JS25] and [ST25] and show how they can be applied to our derandomization of linear codes via the AEL procedure to yield near-liner time list decoding and list recovery algorithms. We start with the list-decoding case.
Theorem 6.15 (Implicit version of Theorem 4.1 from [JS25]).
Let $\mathcal{C}_{\textnormal{in}}$ be a $(\delta_{\textnormal{in}},L_{\textnormal{in}})$-list-decodable code with block length $d$ and $C_{out}$ be uniquely decodable from $\delta_{\textnormal{out}}$ fraction of errors, having block length $N$. Then the AEL code $\mathcal{C}_{\textnormal{AEL}}=C(\mathcal{C}_{\textnormal{in}},\mathcal{C}_{\textnormal{out}},G)$ is $(\rho-\varepsilon,L)$-list-decodable for any constant $0<\varepsilon<\rho$, if $\tfrac{\lambda}{d}<O\left(\tfrac{\varepsilon^{2}\delta_{\textnormal{out}}^{2}}{L_{in}^{4}}\right)$, where $L$ is an upper bound on the list size of $\mathcal{C}_{\textnormal{AEL}}$. Furthermore, if $C_{out}$ is uniquely decodable from $\delta_{\textnormal{out}}$ fraction of errors in time $\mathcal{T}$, then the decoding can be performed in $\widetilde{\mathcal{O}}_{\varepsilon,\ell}(N+\mathcal{T})$ time.
Let $\mathcal{P}$ be an $(L+1)$-LCL property that is the complement of being list-decodable at the $\varepsilon$-relaxed Generalized Singleton Bound. Lemma˜5.31 guarantees that a code $\mathcal{C}_{\textnormal{in}}$ satisfying the complement of $\mathcal{P}$ can be found in time that is constant in the block length $N$. For $C_{out}$, we choose the same code family from [JS25], dating back to [GI05], that has linear time decoders. As a consequence, we obtain near-linear time list-decodable codes up to the $\varepsilon$-relaxed singled codes. Moreover, by Corollary˜6.2, $\mathcal{C}_{\textnormal{AEL}}$ can also satisfy the complement of any number of additional LCL properties.
We now turn our attention to efficient list-recovery.
Theorem 6.16 (Implicit version of Theorem 1.3 from [ST25]).
Let $\mathcal{C}_{\textnormal{in}}$ be a $(\rho,\ell,L_{\textnormal{in}})$-list-recoverable code with block length $d$ and $\mathcal{C}_{\textnormal{out}}$ be a code with block length $N$ that is uniquely decodable from $\delta_{\textnormal{out}}$ fraction of errors. Then the AEL code $\mathcal{C}_{\textnormal{AEL}}=C(\mathcal{C}_{\textnormal{in}},\mathcal{C}_{\textnormal{out}},G)$ is $(\rho-\varepsilon,\ell,L)$-list-recoverable for any constant $0<\varepsilon<\rho$, if $\tfrac{\lambda}{d}<O\left(\tfrac{\varepsilon^{2}\delta_{\textnormal{out}}^{2}}{(\ell\cdot L_{\textnormal{in}})^{2}}\right)$, where $L$ is an upper bound on the output list size of $\mathcal{C}_{\textnormal{AEL}}$. Furthermore, if $\mathcal{C}_{\textnormal{out}}$ is uniquely decodable from $\delta_{\textnormal{out}}$ fraction of errors in time $\mathcal{T}$, then the decoding can be performed in $\widetilde{\mathcal{O}}_{\varepsilon,\ell,L_{in}}(n+\mathcal{T})$ time.
Similar to the above theorem, we can take $\mathcal{C}_{\textnormal{in}}$ to be a code guaranteed by Lemma˜5.31, that achieves the same list-size tradeoffs as random linear codes. The choice of $\mathcal{C}_{\textnormal{out}}$ is the same as above. As a consequence, we obtain near-linear time list-recoverable codes up to the optimal [BCDZ25a] bound. Moreover, by Corollary˜6.2, $\mathcal{C}_{\textnormal{AEL}}$ can also satisfy the complement of any number of additional LCL properties.
@@ 7 Acknowledgements
We thank Jonathan Mosheiff for providing feedback and comments on a draft of the paper, and also thank Yeyuan Chen for providing comments on an earlier version of the paper.